Let x be a normal random variable with mean 20 and standard deviation 4. find p(16<X<22).

A) 0.4672

B) 0.5328

C) 0.1498

D) 0.0328
Answer #1

Hello

YOUR REQUIRED ANSWER IS OPTION B : 0.5328

Given :

X is a normal random variable with \mu = 20 and \sigma = 4.

We can find the z-score using the formula:

z = \frac{x - \mu}{\sigma}

Hence, z score for 16 is :

16-20

and z score for 22 is :

z = \frac{22 - 20}{4} = 0.5

Hence, the required probability p(16<x<22) = p(-1 < z < 0.5) = p(z < 0.5) - p(z < -1)

= 0.691462 - 0.158655 = 0.532807

I hope this solves your query.

Feel free to comment if you still have any query.

Do give a thumbs up if you find this helpful.